OpenStudy (anonymous):

how do I find the vertex and intercepts of 1/3(x^2 +5x-4)?

OpenStudy (muzzack):

Final result : \(\huge\frac{x^2 + 5x - 4}{3}\)

OpenStudy (zzr0ck3r):

\(f(x) = \frac{1}{3}(x^2+5x-4)=\frac{1}{3}(x+\frac{5}{2})^2-4-\frac{1}{3}(\frac{5}{2})^2\) Simplify this and you have vertex form.

OpenStudy (triciaal):

vertex x value is -b/2a and y value is f(x) from the eqn in standard form

OpenStudy (anonymous):

\(\Large \frac{1}{3}(x^2+5x-4)=\frac{x^2}{3} + \frac{5}{3}x-\frac{4}{3}\) The given equation is \(\Large y = \frac{x^2}{3} + \frac{5}{3}x-\frac{4}{3}\) This is of the form \(\Large ax^2 + bx + c = y \) The \(\Large x\) co-ordinate of the vertex \(\Large = \frac{-b}{2a}=\huge \frac{\frac{-5}{3}}{\frac{2}{3}}=\frac{-5}{2}\) Replacing \(\Large x\) by \(\Large \frac{-5}{2}\) in the given equation, \(\Large y=\frac{1}{3}\times (\frac{-5}{2})^2+\frac{5}{3}\times \frac{-5}{2}-\frac{4}{3}=\frac{-41}{12}\) The vertex =\(\huge \color{red}{ (\frac{-5}{2},\frac{-41}{12}) }\) To find the y intercept we replace x by 0. \(\Large \Rightarrow y=\frac{-4}{3}\) To find the x intercept we set y = 0. \(\Large \Rightarrow \frac{x^2}{3} + \frac{5}{3}x-\frac{4}{3}=0\) \(\huge \Rightarrow x=\frac{-5\pm \sqrt{41}}{2}\) Thus the \(\Large x\) intercepts are \(\huge \color{red}{\frac{-5-\sqrt{41}}{2}}\) and \(\huge \color{red}{\frac{-5+\sqrt{41}}{2}}\) while the \(\Large y\) intercept is \(\huge \color{red}{\frac{-4}{3}}\)